OpenStudy (anonymous):

6 Pencils for 13 Dollars Write the rate as a unit rate for the above problem and solve the problem

OpenStudy (texaschic101):

A unit rate has a denominator of 1...you are wanting to find out how much 1 pencil costs. 6/13 = 1/x....6 pencils to $13 = 1 pencil for $x now we cross multiply because this is a proportion 6x = 13 -- now divide by 6 on both sides x = 13/6 x = 2.166 rounds to $2.17 so the unit rate is 2.17/1

OpenStudy (kweenlexis):

yup i agree with @texaschic101

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

:) confirmation is always good

OpenStudy (kweenlexis):

;)

OpenStudy (kweenlexis):

@texaschic101 do you know science?